OpenStudy (anonymous):

determine whether the formula determines y as a function of x if not explain why. 2) x=2y^2

OpenStudy (anonymous):

y^2 = x/2 y = +sqrt (x/2) or - sqrt (x/2) this is not a function because there are two possible outputs so the formula does not determine y as function of x

OpenStudy (anonymous):

what does it mean when they say to determine y as a function of x?

OpenStudy (anonymous):

well find y in terms of x as i have done and check if y is a function of x a one-to-many relation like y = sqrt(x/2) is not a function

OpenStudy (anonymous):

ohh alright i get it now thank you sooooooo much!!!! =]

OpenStudy (anonymous):

yw